A chemist needs 10 liters of a 50% salt solution. all she has available is a 20% salt solution and a 70% salt solution. how much of each of the two solutions should she mix to obtain her desired solution?
Found 2 solutions by richwmiller, nerdybill:
Answer by richwmiller(17219)

a+b=10,
0.2*a+0.7*b=0.5*10
a=10-b
0.2*(10-b)+0.7*b=5
2-0.2b+0.7*b=5
0.5*b=3
b=6
a=10-b
a=4 b=6

Answer by nerdybill(7384)

A chemist needs 10 liters of a 50% salt solution. all she has available is a 20% salt solution and a 70% salt solution. how much of each of the two solutions should she mix to obtain her desired solution?
.
Let x = amount (liters) of 70%
then
10-x = amount (liters) of 20%
.
.70x + .20(10-x) = .50(10)
.70x + 2-.20x = 5
.50x + 2 = 5
.50x = 3
x = 6 liters (70%)
.
amount of 20%:
10-x = 10-6 = 4 liters